Qd = 180 - 2P,

Qs = ‐ 15 + P

The market is government-regulated with a price support per unit and production quotas. If the price is set at $72 per unit, what production quota is needed to make sure there are no shortages or surpluses?


a.an increase in the demand for the good. new demand equation is Qd = 190 - 2P. The government is trying to decide between two options:

Maintain the number of quotas and let the market adjust, or

Maintain the price support and increase the number of quotas.


Suppose that the government decides to maintain the number of quotas and let the market adjust, what is:

The price observed in the market?

The consumer surplus?

The producer surplus?

The deadweight loss?


b. Suppose the government decides to increase the number of quotas available to 72 units but it keeps the price support at the current level of $72, what is:

The consumer surplus?

The producer surplus?

The deadweight loss?


c. Which of the government options in 1a. will be preferred by: The Producers? The Society?


1
Expert's answer
2021-10-14T10:49:01-0400

Solution:

First, derive the equilibrium price and quantity:

At equilibrium: Qd = Qs

180 – 2P = -15 + P

180 + 15 = P + 2P

195 = 3P

P = 65

Equilibrium price = 65

Substitute in either the demand or supply function to derive quantity:

Qd = 180 – 2P = 180 – 2(65) = 180 – 130 = 50

Equilibrium quantity = 50

The price of $72 is above the equilibrium price of $65. When a price floor is set above the equilibrium price, quantity supplied will exceed quantity demanded, and excess supply or surpluses will arise in the market.

The government can, therefore, set a production quota of 30 units to eliminate surpluses in the market.

 

a.). New Qd = 190 – 2P

Calculate new equilibrium price and quantity:

190 – 2P = -15 + P

190 + 15 = P + 2P

205 = 3P

P = 68

New equilibrium price = 68

The government should Maintain the number of quotas and let the market adjust since the price is decreasing.

 

The price observed in the market = $68

The consumer surplus = 45

The producer surplus = 225

The deadweight loss = 30


c.). Price controls will be preferred by the society, while the producers will prefer production quotas.